# Towards Adapting Federated & Quantum Machine Learning for Network Intrusion Detection: A Survey

### v1

@@ I Introduction
With the rapid evolution in modern networks like IoT devices, 5G technologies, and cloud services [1, 2, 3], it has been important to maintain the security of these networks. These networks facilitate huge amounts of sensitive and confidential data sharing, making their security critical. Any compromise to the data can lead to significant consequences. For instance, the lawsuit against Ring highlights the consequences of data leaks, where inadequate security measures allowed hackers to exploit smart home devices, leading to harassment, threats, and severe privacy violations [4]. Network Intrusion Detection System (NIDS) plays a crucial role in safeguarding the network data and identifying any malicious activities within the network [5]. Traditional IDS are trained using a centralized approach, where all the data are transferred to a centralized system for model training. This encounters significant limitations due to large bandwidth requirements for communication, experiencing detection delays and privacy consequences [6]. This limits the IDS’s capability to timely detect evolving cyber risks and highly refined intrusions.
Communication plays an important role in a network of devices, particularly in IoT networks. The interconnectedness of IoT devices is expected to reach 40 billion by 2030 [7]. Identifying anomalous behavior in wireless sensor networks must balance accuracy with energy and communication efficiency [8]. Prior distributed approaches have attempted to address this challenge by reducing communication overhead [9, 10], yet they still required sharing processed data across nodes. Federated Learning (FL) advances this idea by eliminating the need to transfer even processed sensor measurements. Instead, only model parameters are communicated with the server, thereby further reducing communication costs while preserving data privacy compared to traditional centralized methods [11, 12]. Though FL techniques require the frequent transmission of local model updates, various techniques such as model compression [13], asynchronous updates [14] and communication efficient aggregation [15] can be used to mitigate communication bottlenecks. In the centralized method, the model is trained only when all the actual raw data from end devices are transferred to the server. However, IoT devices often operate in remote geographical locations with limited computational power, energy and network connectivity, which can lead to communication failures, preventing some devices from sending their data to the server for training [16].
@@
To address these challenges, the FL approach has been introduced [17]. FL can be used to produce a global model that is dynamically updated. Fig. 1 shows a high level overview of FL in IDS. When an end device experiences a loss of network or energy, it can defer transmitting its local model updates. Once the device regains connectivity and sufficient power, it can send its updates which are then dynamically incorporated into the global model [18]. The global model [19] produced by FL synthesizes knowledge from all the participating clients, capturing insights and patterns from their diverse data while maintaining data privacy.
There are various challenges in analyzing the data that are generated across multiple distributed nodes (data sources) [20], including:
@@
Traditional machine learning models, such as support vector machine (SVM), K-nearest neighbor (KNN), naïve Bayes, logistic regression (LR), decision trees, clustering, and combined and hybrid methods have been widely used in NIDS. However, machine learning (ML) models often rely on handcrafted features, requiring domain expertise and extensive pre-processing to ensure optimal performance [21]. This manual feature engineering can limit the scalability and adaptability of traditional ML approaches in dynamic and complex network environments. Deep Learning (DL), on the other hand, has shown great accuracy in detecting sophisticated network attacks. It automates feature extraction and leverages hierarchical representations of data. DL models can generalize better in heterogeneous network conditions and dynamically adapt to evolving attack patterns [21]. This makes them a powerful alternative to traditional ML approaches for NIDS, especially in environments where accuracy, scalability, and minimal human intervention are critical. While DL often requires more computational resources than ML, advancements in hardware and distributed frameworks like FL have mitigated these challenges, further enhancing the viability of deep learning for modern NIDS.
The research in QML is gaining popularity as it can be used to handle high-dimensional complex network traffic data [22]. Although DL methods can detect intricate patterns, they often struggle as the volume and dimensionality of the data increase. QML, leveraging quantum phenomena such as superposition and entanglement, offers a new computational paradigm that could provide significant advantages in representational power and learning efficiency. In particular, researchers are exploring hybrid quantum-classical methods that could provide significant advantages in representing high-dimensional complex network data and increasing learning efficiency. However, in the real world, the practical deployment of QML is currently constrained due to limited access to actual quantum hardware. Despite this, early research indicates that QML may hold the key to building more efficient and robust IDS in the future.
@@ II-A Intrusion Detection System (IDS)
In the realm of network security, ensuring the confidentiality, integrity, and availability of data is important. IDS are critical in safeguarding networks and systems against malicious activities and unauthorized access [23]. An IDS is a security solution that monitors and analyzes network traffic, system logs, and activities to detect potential threats, such as unauthorized access, malware, and policy violations.
IDS can be broadly classified, based on their monitoring strategies and deployment, into the following categories [6, 23]:
@@
NIDS are classified into two categories based on their detection methodology: signature-based detection and anomaly-based detection. Traditional IDS are based on signature-based anomaly detection and they use the traditional machine learning algorithms to detect anomalies in network traffic. Signature-based detection relies on predefined patterns, known as signatures, of known attacks. These signatures are essentially fingerprints of malicious activities, derived from historical data and prior analysis [24]. The system matches incoming network traffic against this signatures database to identify intrusions. Though this traditional method can detect known attacks/anomalies with higher accuracy, it is inefficient in detecting unknown anomalies or zero-day attacks [23]. As network data continues to evolve with emerging threats and novel attack patterns, there is a critical need for an alternative approach to address these challenges. This is where anomaly-based detection becomes essential, providing a method to identify deviations from normal behavior that may indicate potential attacks [24]. This method leverages statistical models, machine learning, or deep learning to characterize normal traffic and flag any deviations as potential intrusions. It is very efficient in detecting zero-day attacks. The main disadvantage of this method is that it can lead to high false positives [23]. It also requires large volumes of normal behavior data for training, which can be challenging to obtain. Combining signature and anomaly-based techniques, hybrid systems aim to achieve a balance between accuracy and adaptability, improving detection rates while minimizing false alarms.
@@ II-B Deep Learning in NIDS
NIDS are essential tools in cyber security, that are used to identify and respond to malicious activities within a network [5]. Traditional NIDS techniques, such as signature-based and rule-based methods, have limitations in detecting unknown or evolving attack patterns. To address these limitations, deep learning (DL), a powerful subset of machine learning, has gained significant attention for enhancing NIDS performance. Deep learning models like Convolutional Neural Networks(CNNs), Recurrent Neural Networks(RNNs) and Autoencoders have shown promising results in NIDS [25, 26, 27]. CNNs are effective at identifying spatial patterns in network traffic [28]. RNNs are developed to process sequential data to capture temporal attack behaviors [29], and Autoencoders are particularly useful for anomaly detection by learning the normal patterns of network activity [25].
@@
Recent advancements in deep learning for NIDS have explored hybrid models, combining the strengths of traditional machine learning with deep learning techniques [30], as well as innovations like reinforcement learning, model compression techniques and federated learning [31], which offer solutions to data privacy and resource constraints.
@@ II-C1 Concept
The main objective of federated learning is to achieve performance that is as close as possible to centralized learning while addressing the challenges of data privacy and security [32]. In mathematical terms, this means minimizing the difference between the loss function of the global model in both approaches. The objective can be expressed as:
@@ II-C2 Federated Learning Life cycle
The federated learning life cycle consists of 6 stages [33] (see Fig. 2(b)) as follows:
@@
    Model Aggregation and Update: The server aggregates all the received local updates using a predefined method such as FedAvg (Federated Averaging) [34], and this combined update is used to adjust the global model.
@@ III Related Works
In this section, we discuss existing research papers on FL and explain how this survey paper advances beyond them. Yang et al. [32] summarized FL, covering the foundational concepts and applications of FL. The paper discusses FL architectures, privacy preservation techniques, and future directions in the field. However, it lacks a detailed discussion on deployment scenarios, communication methods, and aggregation strategies, which are crucial for the practical implementation of FL in real-world scenarios.
@@
The interconnectedness of IoT networks raises significant security concerns due to the sensitive data recorded by various edge nodes. Aledhari et al. [35] summarized FL technologies, protocols, and applications. The paper delves deeply into FL architectures, various frameworks, network protocols, and the applications of FL, including its use in IDS. However, it lacks in-depth discussion on various aggregation techniques and deployment scenarios and advanced privacy-preserving techniques, which are vital for practical and secure implementations of FL in real-world scenarios. Zhang et al. [20] presented a comprehensive survey on FL. The paper introduced the existing work of federated learning from five aspects: Data Partitioning, Privacy Mechanism, Machine Learning Model, Communication Architecture, and Systems Heterogeneity. The survey offers an in-depth analysis of these aspects, providing foundational insights into the technical and theoretical aspects of FL. However, the paper has notable limitations. It does not explore the applicability of FL in IDS, a critical area where privacy-preserving methods could have a significant impact given the sensitivity of network traffic data. Additionally, the survey lacks a detailed discussion on FL deployment scenarios, a crucial aspect for practical implementation in real-world systems. Another gap is the limited coverage of diverse model aggregation techniques, which plays a central role in achieving effective global models while addressing challenges such as communication overhead and heterogeneity among participating clients. Nguyen et al. [36] reviewed the application of FL in IoT networks, emphasizing how FL can preserve the privacy of sensitive data while enabling collaborative learning. The survey extensively discusses the applications of FL in IoT, its network structures, and architectures, providing valuable insights into the intersection of FL and IoT. However, the survey lacks a detailed exploration of deployment scenarios, which are critical for real-world implementation, as well as an in-depth analysis of model aggregation techniques that address the challenges of data heterogeneity and communication efficiency. Additionally, the absence of a discussion on standardized frameworks for implementing FL in IoT networks limits its utility for practitioners aiming for scalable and robust deployments. Mothukuri et al. [37] presented an overview of the security and privacy aspects of FL. The paper provides an in-depth overview of FL, including various aggregation techniques and architectures, and a detailed classification of security threats and attacks within the FL domain. Additionally, it extensively discusses defensive techniques to mitigate these vulnerabilities, making it a valuable resource for researchers exploring security-centric FL methodologies. However, the paper is primarily focused on security and privacy concerns and does not cover network protocols in detail, which are essential for addressing communication challenges in FL systems.
@@
In a world where data privacy and collaborative intelligence converge, FL emerges as a transformative solution. Abdulrahman et al. [38] discussed the transition from centralized approaches to FL. The paper provides a detailed discussion of the privacy and security concerns and challenges associated with FL, its applications, and the optimization of federated algorithms. However, the paper lacks a focused survey on the use of FL for IDS and an exploration of deployment scenarios in practical applications. It also does not mention deep learning using FL. Li et al. [39] discussed FL, providing an overview of its fundamental concepts, including various FL architectures, different machine learning models in FL, deployment scenarios, and communication architectures. However, the paper lacks a detailed discussion on various aggregation techniques and the application of FL in IDS. Banabilah et al. [40] presented a survey on FL, discussing various FL architectures, their applications, and the challenges associated with their implementation. However, the paper lacks a detailed discussion on deployment strategies, various aggregation algorithms, and the specific applications of FL in IDS. Wen et al. [41] presented a survey on FL that provides foundational knowledge on the subject. The paper addresses key aspects of FL, including privacy and security protection mechanisms, communication overhead challenges, and heterogeneity issues. Furthermore, it highlights various applications of FL, including its use in IDS. However, the survey falls short in several critical areas. It lacks detailed discussions on standardized frameworks, effective communication techniques, and practical deployment scenarios for FL. Moreover, the paper does not provide an in-depth analysis of the diverse model aggregation techniques crucial for optimizing global models, particularly in environments characterized by high data heterogeneity and constrained resources. Khraisat et al. [42] provided a comprehensive overview of FL, addressing key aspects such as FL architectures, communication methods, aggregation techniques, and FL frameworks. However, the study does not delve into federated deep learning specifically for IDS and lacks coverage of network protocols.
@@
Feature Yang et al. (2019) [32] Aledhari et al. (2020) [35] Zhang et al. (2021) [20] Nguyen et al. (2021) [36] Mothukuri et al. (2021) [37] Abdulrahman et al. (2021) [38] Li et al. (2021) [39] Banabilah et al. (2022) [40] Wen et al. (2023) [41] Khraisat et al. (2024) [42] This paper
@@ IV-A Types of Federated Learning
FL can be categorized into various types depending on how the data is distributed and how the model is trained across the participants. The primary types of FL are Horizontal Federated Learning (HFL), Vertical Federated Learning (VFL), and Federated Transfer Learning (FTL) (see Fig. 4 )[43].
Horizontal Federated Learning (HFL): In Horizontal Federated Learning, the clients(devices) have the same set of features (attributes) but different samples (data instances)(see Fig. 4(a)). This is typically the case when each client has data that represents a subset of the overall population, but the data is homogeneous in terms of features. Example: A smartphone app that collects health data, such as heart rate and steps taken. Different users’ phones generate records with the same set of features but different data instances [44].
Vertical Federated Learning (VFL): In Vertical Federated Learning, the clients have different features(attributes) for the same set of samples(data instances)(see Fig. 4(b)). This is typically the case when multiple clients have data about the same entities but with different attributes or features. Example: Two hospitals may have data on the same patients(same data instances), but one hospital has medical history data(e.g., diagnoses, treatments), while the other has lab results(e.g., blood tests, imaging reports) [45].
Federated Transfer Learning (FTL): Federated Transfer Learning (FTL) combines the principles of both Federated Learning and Transfer Learning (see Fig. 4(c)). In FTL, a model is trained across a set of clients with limited data or resources in such a way that knowledge from related domains or tasks is transferred to improve the learning process in a target domain. It is especially useful in situations where one client has insufficient data to train a model effectively but can leverage knowledge from other clients or previously trained models. Example: A model trained on one set of medical data(e.g., detecting pneumonia from X-ray images) can be fine-tuned with a smaller dataset from a different hospital to detect other types of medical conditions, such as tuberculosis, without requiring extensive data from the new hospital [46].

### v2

@@ I Introduction
With the rapid evolution in modern networks like IoT devices, 5G technologies, and cloud services [10112761, dangi2021study, rao2021evolving], it has been important to maintain the security of these networks. These networks facilitate huge amounts of sensitive and confidential data sharing, making their security critical. Any compromise to the data can lead to significant consequences. For instance, the lawsuit against Ring highlights the consequences of data leaks, where inadequate security measures allowed hackers to exploit smart home devices, leading to harassment, threats, and severe privacy violations [theguardian2020amazon]. Network Intrusion Detection System (NIDS) plays a crucial role in safeguarding the network data and identifying any malicious activities within the network [ahmad2021network]. Traditional IDS are trained using a centralized approach, where all the data are transferred to a centralized system for model training. This encounters significant limitations due to large bandwidth requirements for communication, experiencing detection delays and privacy consequences [khraisat2019survey]. This limits the IDS’s capability to timely detect evolving cyber risks and highly refined intrusions.
Communication plays an important role in a network of devices, particularly in IoT networks. The interconnectedness of IoT devices is expected to reach 40 billion by 2030 [iot_devices]. Identifying anomalous behavior in wireless sensor networks must balance accuracy with energy and communication efficiency [rajasegarar2008anomaly]. Prior distributed approaches have attempted to address this challenge by reducing communication overhead [rajasegarar2006distributed, rajasegarar2014ellipsoidal], yet they still required sharing processed data across nodes. Federated Learning (FL) advances this idea by eliminating the need to transfer even processed sensor measurements. Instead, only model parameters are communicated with the server, thereby further reducing communication costs while preserving data privacy compared to traditional centralized methods [zhou2021communication, dritsas2025federated]. Though FL techniques require the frequent transmission of local model updates, various techniques such as model compression [le2024survey], asynchronous updates [wang2022asynchronous] and communication efficient aggregation [bonawitz2019federated] can be used to mitigate communication bottlenecks. In the centralized method, the model is trained only when all the actual raw data from end devices are transferred to the server. However, IoT devices often operate in remote geographical locations with limited computational power, energy and network connectivity, which can lead to communication failures, preventing some devices from sending their data to the server for training [o2014anomaly].
@@
To address these challenges, the FL approach has been introduced [li2020review]. FL can be used to produce a global model that is dynamically updated. Fig. 1 shows a high level overview of FL in IDS. When an end device experiences a loss of network or energy, it can defer transmitting its local model updates. Once the device regains connectivity and sufficient power, it can send its updates which are then dynamically incorporated into the global model [mughal2024adaptive]. The global model [qi2024model] produced by FL synthesizes knowledge from all the participating clients, capturing insights and patterns from their diverse data while maintaining data privacy.
There are various challenges in analyzing the data that are generated across multiple distributed nodes (data sources) [zhang2021survey], including:
@@
Traditional machine learning models, such as support vector machine (SVM), K-nearest neighbor (KNN), naïve Bayes, logistic regression (LR), decision trees, clustering, and combined and hybrid methods have been widely used in NIDS. However, machine learning (ML) models often rely on handcrafted features, requiring domain expertise and extensive pre-processing to ensure optimal performance [liu2019machine]. This manual feature engineering can limit the scalability and adaptability of traditional ML approaches in dynamic and complex network environments. Deep Learning (DL), on the other hand, has shown great accuracy in detecting sophisticated network attacks. It automates feature extraction and leverages hierarchical representations of data. DL models can generalize better in heterogeneous network conditions and dynamically adapt to evolving attack patterns [liu2019machine]. This makes them a powerful alternative to traditional ML approaches for NIDS, especially in environments where accuracy, scalability, and minimal human intervention are critical. While DL often requires more computational resources than ML, advancements in hardware and distributed frameworks like FL have mitigated these challenges, further enhancing the viability of deep learning for modern NIDS.
The research in QML is gaining popularity as it can be used to handle high-dimensional complex network traffic data [amin2018quantum]. Although DL methods can detect intricate patterns, they often struggle as the volume and dimensionality of the data increase. QML, leveraging quantum phenomena such as superposition and entanglement, offers a new computational paradigm that could provide significant advantages in representational power and learning efficiency. In particular, researchers are exploring hybrid quantum-classical methods that could provide significant advantages in representing high-dimensional complex network data and increasing learning efficiency. However, in the real world, the practical deployment of QML is currently constrained due to limited access to actual quantum hardware. Despite this, early research indicates that QML may hold the key to building more efficient and robust IDS in the future.
@@ II-A Intrusion Detection System (IDS)
In the realm of network security, ensuring the confidentiality, integrity, and availability of data is important. IDS are critical in safeguarding networks and systems against malicious activities and unauthorized access [liao2013intrusion]. An IDS is a security solution that monitors and analyzes network traffic, system logs, and activities to detect potential threats, such as unauthorized access, malware, and policy violations.
IDS can be broadly classified, based on their monitoring strategies and deployment, into the following categories [khraisat2019survey, liao2013intrusion]:
@@
NIDS are classified into two categories based on their detection methodology: signature-based detection and anomaly-based detection. Traditional IDS are based on signature-based anomaly detection and they use the traditional machine learning algorithms to detect anomalies in network traffic. Signature-based detection relies on predefined patterns, known as signatures, of known attacks. These signatures are essentially fingerprints of malicious activities, derived from historical data and prior analysis [alazab2021federated]. The system matches incoming network traffic against this signatures database to identify intrusions. Though this traditional method can detect known attacks/anomalies with higher accuracy, it is inefficient in detecting unknown anomalies or zero-day attacks [liao2013intrusion]. As network data continues to evolve with emerging threats and novel attack patterns, there is a critical need for an alternative approach to address these challenges. This is where anomaly-based detection becomes essential, providing a method to identify deviations from normal behavior that may indicate potential attacks [alazab2021federated]. This method leverages statistical models, machine learning, or deep learning to characterize normal traffic and flag any deviations as potential intrusions. It is very efficient in detecting zero-day attacks. The main disadvantage of this method is that it can lead to high false positives [liao2013intrusion]. It also requires large volumes of normal behavior data for training, which can be challenging to obtain. Combining signature and anomaly-based techniques, hybrid systems aim to achieve a balance between accuracy and adaptability, improving detection rates while minimizing false alarms.
@@ II-B Deep Learning in NIDS
NIDS are essential tools in cyber security, that are used to identify and respond to malicious activities within a network [ahmad2021network]. Traditional NIDS techniques, such as signature-based and rule-based methods, have limitations in detecting unknown or evolving attack patterns. To address these limitations, deep learning (DL), a powerful subset of machine learning, has gained significant attention for enhancing NIDS performance. Deep learning models like Convolutional Neural Networks(CNNs), Recurrent Neural Networks(RNNs) and Autoencoders have shown promising results in NIDS [chen2018autoencoder, mohammadpour2022survey, tang2018deep]. CNNs are effective at identifying spatial patterns in network traffic [kim2020cnn]. RNNs are developed to process sequential data to capture temporal attack behaviors [al2007power], and Autoencoders are particularly useful for anomaly detection by learning the normal patterns of network activity [chen2018autoencoder].
@@
Recent advancements in deep learning for NIDS have explored hybrid models, combining the strengths of traditional machine learning with deep learning techniques [sajid2024enhancing], as well as innovations like reinforcement learning, model compression techniques and federated learning [agrawal2022federated], which offer solutions to data privacy and resource constraints.
@@ II-C1 Concept
The main objective of federated learning is to achieve performance that is as close as possible to centralized learning while addressing the challenges of data privacy and security [yang2019federated]. In mathematical terms, this means minimizing the difference between the loss function of the global model in both approaches. The objective can be expressed as:
@@ II-C2 Federated Learning Life cycle
The federated learning life cycle consists of 6 stages [li2024threats] (see Fig. 2(b)) as follows:
@@
    Model Aggregation and Update: The server aggregates all the received local updates using a predefined method such as FedAvg (Federated Averaging) [li2019convergence], and this combined update is used to adjust the global model.
@@ III Related Works
In this section, we discuss existing research papers on FL and explain how this survey paper advances beyond them. Yang et al. [yang2019federated] summarized FL, covering the foundational concepts and applications of FL. The paper discusses FL architectures, privacy preservation techniques, and future directions in the field. However, it lacks a detailed discussion on deployment scenarios, communication methods, and aggregation strategies, which are crucial for the practical implementation of FL in real-world scenarios.
@@
The interconnectedness of IoT networks raises significant security concerns due to the sensitive data recorded by various edge nodes. Aledhari et al. [9153560] summarized FL technologies, protocols, and applications. The paper delves deeply into FL architectures, various frameworks, network protocols, and the applications of FL, including its use in IDS. However, it lacks in-depth discussion on various aggregation techniques and deployment scenarios and advanced privacy-preserving techniques, which are vital for practical and secure implementations of FL in real-world scenarios. Zhang et al. [zhang2021survey] presented a comprehensive survey on FL. The paper introduced the existing work of federated learning from five aspects: Data Partitioning, Privacy Mechanism, Machine Learning Model, Communication Architecture, and Systems Heterogeneity. The survey offers an in-depth analysis of these aspects, providing foundational insights into the technical and theoretical aspects of FL. However, the paper has notable limitations. It does not explore the applicability of FL in IDS, a critical area where privacy-preserving methods could have a significant impact given the sensitivity of network traffic data. Additionally, the survey lacks a detailed discussion on FL deployment scenarios, a crucial aspect for practical implementation in real-world systems. Another gap is the limited coverage of diverse model aggregation techniques, which plays a central role in achieving effective global models while addressing challenges such as communication overhead and heterogeneity among participating clients. Nguyen et al. [nguyen2021federated] reviewed the application of FL in IoT networks, emphasizing how FL can preserve the privacy of sensitive data while enabling collaborative learning. The survey extensively discusses the applications of FL in IoT, its network structures, and architectures, providing valuable insights into the intersection of FL and IoT. However, the survey lacks a detailed exploration of deployment scenarios, which are critical for real-world implementation, as well as an in-depth analysis of model aggregation techniques that address the challenges of data heterogeneity and communication efficiency. Additionally, the absence of a discussion on standardized frameworks for implementing FL in IoT networks limits its utility for practitioners aiming for scalable and robust deployments. Mothukuri et al. [mothukuri2021survey] presented an overview of the security and privacy aspects of FL. The paper provides an in-depth overview of FL, including various aggregation techniques and architectures, and a detailed classification of security threats and attacks within the FL domain. Additionally, it extensively discusses defensive techniques to mitigate these vulnerabilities, making it a valuable resource for researchers exploring security-centric FL methodologies. However, the paper is primarily focused on security and privacy concerns and does not cover network protocols in detail, which are essential for addressing communication challenges in FL systems.
@@
In a world where data privacy and collaborative intelligence converge, FL emerges as a transformative solution. Abdulrahman et al. [9220780] discussed the transition from centralized approaches to FL. The paper provides a detailed discussion of the privacy and security concerns and challenges associated with FL, its applications, and the optimization of federated algorithms. However, the paper lacks a focused survey on the use of FL for IDS and an exploration of deployment scenarios in practical applications. It also does not mention deep learning using FL. Li et al. [li2021survey] discussed FL, providing an overview of its fundamental concepts, including various FL architectures, different machine learning models in FL, deployment scenarios, and communication architectures. However, the paper lacks a detailed discussion on various aggregation techniques and the application of FL in IDS. Banabilah et al. [banabilah2022federated] presented a survey on FL, discussing various FL architectures, their applications, and the challenges associated with their implementation. However, the paper lacks a detailed discussion on deployment strategies, various aggregation algorithms, and the specific applications of FL in IDS. Wen et al. [wen2023survey] presented a survey on FL that provides foundational knowledge on the subject. The paper addresses key aspects of FL, including privacy and security protection mechanisms, communication overhead challenges, and heterogeneity issues. Furthermore, it highlights various applications of FL, including its use in IDS. However, the survey falls short in several critical areas. It lacks detailed discussions on standardized frameworks, effective communication techniques, and practical deployment scenarios for FL. Moreover, the paper does not provide an in-depth analysis of the diverse model aggregation techniques crucial for optimizing global models, particularly in environments characterized by high data heterogeneity and constrained resources. Khraisat et al. [khraisat2024survey] provided a comprehensive overview of FL, addressing key aspects such as FL architectures, communication methods, aggregation techniques, and FL frameworks. However, the study does not delve into federated deep learning specifically for IDS and lacks coverage of network protocols.
@@
Feature Yang et al. (2019) [yang2019federated] Aledhari et al. (2020) [9153560] Zhang et al. (2021) [zhang2021survey] Nguyen et al. (2021) [nguyen2021federated] Mothukuri et al. (2021) [mothukuri2021survey] Abdulrahman et al. (2021) [9220780] Li et al. (2021) [li2021survey] Banabilah et al. (2022) [banabilah2022federated] Wen et al. (2023) [wen2023survey] Khraisat et al. (2024) [khraisat2024survey] This paper
@@ IV-A Types of Federated Learning
FL can be categorized into various types depending on how the data is distributed and how the model is trained across the participants. The primary types of FL are Horizontal Federated Learning (HFL), Vertical Federated Learning (VFL), and Federated Transfer Learning (FTL) (see Fig. 4 )[10.1145/3511285.3511291].
Horizontal Federated Learning (HFL): In Horizontal Federated Learning, the clients(devices) have the same set of features (attributes) but different samples (data instances)(see Fig. 4(a)). This is typically the case when each client has data that represents a subset of the overall population, but the data is homogeneous in terms of features. Example: A smartphone app that collects health data, such as heart rate and steps taken. Different users’ phones generate records with the same set of features but different data instances [nguyen2022federated].
Vertical Federated Learning (VFL): In Vertical Federated Learning, the clients have different features(attributes) for the same set of samples(data instances)(see Fig. 4(b)). This is typically the case when multiple clients have data about the same entities but with different attributes or features. Example: Two hospitals may have data on the same patients(same data instances), but one hospital has medical history data(e.g., diagnoses, treatments), while the other has lab results(e.g., blood tests, imaging reports) [liu2024vertical].
Federated Transfer Learning (FTL): Federated Transfer Learning (FTL) combines the principles of both Federated Learning and Transfer Learning (see Fig. 4(c)). In FTL, a model is trained across a set of clients with limited data or resources in such a way that knowledge from related domains or tasks is transferred to improve the learning process in a target domain. It is especially useful in situations where one client has insufficient data to train a model effectively but can leverage knowledge from other clients or previously trained models. Example: A model trained on one set of medical data(e.g., detecting pneumonia from X-ray images) can be fine-tuned with a smaller dataset from a different hospital to detect other types of medical conditions, such as tuberculosis, without requiring extensive data from the new hospital [chen2020fedhealth].